OpenStudy (selena2345):

A store marks up items to make a 30% profit. If an item cost $7.50 from the supplier,what will the selling price be ?

OpenStudy (anonymous):

I will help you again ;)

OpenStudy (anonymous):

Ok, to solve, you have to multiply 7.50 by 30% then add that to 7.50

OpenStudy (selena2345):

7.50x30%=2.25+7.50=9.75

OpenStudy (anonymous):

Now add that to 7.50 :)

OpenStudy (selena2345):

i did

OpenStudy (anonymous):

Oh yeah, but that is about it :)

OpenStudy (selena2345):

2.25+7.50=9.75

OpenStudy (anonymous):

That will be your answer :)